OpenStudy (anonymous):

I need help simplifying this power series. (x^2-1)sum_{n=2}^{infty}[n(n-1)C_nx^{n-2}]+4x\sum_{n=1}^{infty}[nC_nx^{n-1}]+2\sum_{n=0}^{infty}[C_nx^{n}] I need to get everything under 1 summation and so that \[x^{n}\] pulls out but I cannot figure out how to do it. Thanks

OpenStudy (anonymous):

\[(x^2-1)\sum_{n=2}^{infty}[n(n-1)C_nx^{n-2}]+4x\sum_{n=1}^{infty}[nC_nx^{n-1}]+2\sum_{n=0}^{infty}[C_nx^{n}]\]

OpenStudy (anonymous):

This is for Diff Equ

OpenStudy (anonymous):

Yeah, im not looking for x really. You use the fact that \[\sum_{i}^{\infty}[K_nx^n]=0 \] means that K_n = 0 and find y (which is where we get these power series) in terms of a new power series.

OpenStudy (loser66):

@satellite73